A diameter is the segment from (-6,-4) to (3,5) ... Find the equation of the circle.
Answer by lwsshak3(11628)

A diameter is the segment from (-6,-4) to (3,5) ... Find the equation of the circle.
**
Standard form of equation for a circle: (x-h)^2+(y-k)^2=r^2, (h,k)=(x,y) coordinates of center, r=radius
Finding mid points of x and y coordinates of given line segment end points:
x=(3-6)/2=-3/2=-1.5
y=(5-4)/2=12=.5
center of circle: (-1.5,0.5)
..
Finding diameter of circle using distance formula:
d^2=(3+6)^2+(5+4)^2=9^2+9^2=81+81=162
d=√162
radius=d/2=√162/2≈6.36
radius^2=40.5
Equation of circle:
(x+1.5)^2+(y-0.5)^2=40.5